# Twisted Superfields

### v8

@@ 3.3 Quark masses and mixing
Most of the off-diagonal FCNC elements of these matrices are small, but a few are more significant. The only Z mixing matrix elements with magnitudes greater than 0.006 are $V_{Zs\bar{d}}^{A},V_{Zf\bar{b}}^{L},V_{Zf\bar{s}}^{L},V_{Zf\bar{d}}^{L},V_{Zb\bar{d}}^{L}$, $V_{Zt\bar{c}}^{R}$, and $V_{Zt\bar{u}}^{R}$. More detailed work would be needed to verify that the $Z$ boson FCNC mixing of this model is fully consistent with experimental data.
3.4 Leptons
In this model, charged leptons get their masses not from the vev of the Higgs boson but from vevs of the $\varphi_{2}$ adjoint scalars (see eq (3.4)). Through the superpotential couplings of eq (1.18), $m_{1FF^{\prime}}$ and $\Gamma_{1FF^{\prime}}$ generate charged lepton masses. Combining those with quantum-generated masses from (A.12) leads to the following lepton mass terms:
$\left(\begin{array}[]{c}{\tilde{\nu}_{W}^{(F)}}\\ {\tilde{e}_{W}^{(F)}}\\ {\tilde{e}^{(F)}}\end{array}\right)^{T}\left(\begin{array}[]{ccc}{\tilde{m}_{1FF^{\prime}}}&{0}&{0}\\ {0}&{\tilde{m}^{\prime}_{1FF^{\prime}}}&{\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}}\\ {0}&{c\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}}&{\tilde{m}^{\prime}_{1FF^{\prime}}}\end{array}\right)\left(\begin{array}[]{c}{\nu_{W}^{(F^{\prime})}}\\ {e_{W}^{(F^{\prime})}}\\ {e^{(F^{\prime})}}\end{array}\right)\,.$ (3.27)
As discussed in the Appendix, quantum effects can generate differences between $\tilde{m}_{1FF^{\prime}}$ and $\tilde{m}^{\prime}_{1FF^{\prime}}$. The $\tilde{m}^{\prime}_{1FF^{\prime}}$ terms are much smaller than the $\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}$ terms.
In that case, since $c\ll 1$, the lepton mass terms for $\tilde{e}_{W}^{(F)}e^{(F^{\prime})}$ are much larger than those for $\tilde{e}^{(F)}e_{W}^{(F^{\prime})}$. For that reason, $\tilde{e}_{W}^{(F)}$ and $e^{(F^{\prime})}$ are mapped to the right- and left-handed components of proposed heavy charged leptons, denoted “Omega” leptons. $\tilde{e}^{(F)}$ and $e_{W}^{(F^{\prime})}$ are mapped to the right- and left-handed components of the known light charged leptons (electron, muon, and tau). The masses of the Omega leptons should be between 103 GeV and $\sim$30-90 TeV. The lower limit comes from the 95% confidence exclusion limit from heavy lepton searches [22]. The upper limit comes from the anomaly-based analysis surrounding eq (4.1).
@@ 3.4 Leptons
From eq (2.18), the last three diagonal components of the photon field when acting on leptons are $e\left(0,1,1\right)$. As a result, if the left- and right-chiral components of the normal light charged leptons ($l$) are mapped to $\bar{l}_{L}^{(F)}=e_{W}^{(F)}$ and $l_{R}^{(F)}=\tilde{e}^{(F)}$, then those leptons have the correct electric charges of $+1$ and $-1$ as well as the correct isodoublet and isosinglet designations. The three flavors of heavy and light leptons ($\Omega$ and $l$) can therefore be identified as follows:
$\begin{array}[]{ccc}{\nu_{W}^{(F)}=\bar{\nu}_{l}^{(F)}}&{e_{W}^{(F)}=\bar{l}_{L}^{(F)}}&{e^{(R)}=\Omega_{L}^{(F)}}\\ {\tilde{\nu}_{W}^{(F)}=\nu_{\Omega}^{(F)}}&{\tilde{e}_{W}^{(F)}=\Omega_{R}^{(F)}}&{\tilde{e}^{(F)}=l_{R}^{(F)},}\end{array}$ (3.28)
@@
From this identification, it can be seen that the negatively charged light leptons have the correct behavior that only their left-handed components interact with the $W$ boson, forming isodoublets with their neutrinos. On the other hand, the opposite is true for negatively charged Omega leptons; their $W$ boson interactions are right-handed. The light leptons not only have the same interactions with the photon and $W$ boson as in the Standard Model, they also have the same interactions with the $Z$ boson, as can be seen from eqs (2.19) and (2.20).
Appendix eq (A.14) shows how quantum effects generate a large Majorana mass $M_{\nu}$ for all three of the $\nu_{\Omega}^{(F)}$ right-handed neutrinos. Again, for reasons discussed around eq (4.1), $M_{\nu}$ should be less than $\sim$30-90 TeV. If $M_{\nu}$ is near that upper limit and the Dirac neutrino mass terms $\tilde{m}_{1FF^{\prime}}$ are around the 1 MeV range, then a seesaw with $M_{\nu}$ will lead to very small masses for the left-handed neutrinos that are in the correct range to match observations. The parameters $\tilde{m}_{1FF^{\prime}}$ may be chosen so that the eigenvalues of the seesaw-generated light neutrino mass matrix reproduce the observed mass differences [23]. For each choice of the nine parameters $\tilde{m}_{1FF^{\prime}}$, diagonalization of the neutrino mass matrix determines the $3\times 3$ left-handed and $3\times 3$ right-handed rotation matrices that are required to transform from the flavor basis of this paper to the neutrino mass eigenbasis.
The masses of the charged Omega leptons are defined by the $3\times 3$ mass matrix $\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}$. The light charged lepton $3\times 3$ mass matrix is partly generated by $c\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}$ and partly by a seesaw between $\bar{\varphi}_{2}\Gamma_{1FF^{\prime}}$ and the $\tilde{m}^{\prime}_{1FF^{\prime}}$ masses. The parameters $\Gamma_{1FF^{\prime}}$, $\tilde{m}^{\prime}_{1FF^{\prime}}$ and $c$ may be chosen so that the eigenvalues of the light charged lepton mass matrix are equal to the electron, muon and tau masses. For each choice of these parameters, diagonalization of the light charged lepton mass matrix determines the $3\times 3$ left-handed and $3\times 3$ right-handed rotation matrices that are required to transform from the flavor basis of this paper to the light charged lepton mass eigenbasis. In particular, these parameters may be chosen so that the combined left-handed rotation from the light neutrino mass eigenbasis to the light charged lepton mass eigenbasis is able to reproduce the observed Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix [23].
@@ 4 Unification and Anomalies
As mentioned in section 1, at a scale above any symmetry breaking, the model of this paper is free of anomalies. The reason is simple: for every fermion in every representation, there is another fermion in a conjugate representation of each of the SU(3) groups that also has opposite charges for the Abelian fields $A_{+\mu}^{0}$ and $A_{-\mu}^{0}$. Consequently, all of the gauge anomaly triangle diagrams cancel.
The unification scale $\Lambda_{U}$ of this model is defined as the scale where the vevs $\bar{\phi}_{23}$ and $\bar{\tilde{\phi}}_{23}$ break the original SU(3)$\times$×SU(3)$\times$×U(1)$\times$×U(1) symmetry down to SU(3)$\times$×SU(2)$\times$×U(1)$\times$×U(1). Below the unification scale, the four couplings $g_{3}$, $g_{2}$, $g_{Y}$ and $g_{Y^{\prime}}$ for these four groups run differently. As mentioned in section 3, this symmetry breaking causes the quark pairs $\tilde{d}^{G}d^{\left(3\right)}$ and $\tilde{d}^{\left(3\right)}d^{G}$ to acquire unification-scale masses. These heavy fermions are in conjugate representations of the remaining SU(3) group, are SU(2) singlets, and have equal and opposite charges for the Abelian $Y$ and $Y^{\prime}$ fields. Consequently, the effective theory below the unification scale that ignores these heavy quarks is also free of local gauge anomalies.
As can be seen from $T_{21}^{Y}$ and $T_{21}^{Y^{\prime}}$ of eqs (2.14) and (2.15), the coupling of right-handed top quarks to $Y^{\prime}$ is $\sqrt{6}$ times larger than to $Y$. The $Y^{\prime}$ charges of all of the other fermions and scalars are also larger than their $Y$ charges. The net result is that below the unification scale, $g_{Y^{\prime}}$ drops to lower scales much more quickly than $g_{Y}$. This is a partial justification for the small value of $\sin\phi_{Z}$ assumed in sections 3 and 5.
Below the unification scale, the next lower scale of the theory is defined by the large Majorana mass $\tilde{M}_{\nu}$ of the three right-handed neutrinos $\tilde{\nu}_{W}^{(F)}$ discussed in eq (A.14) of the Appendix. As a default model, it is assumed that two of the charged Omega leptons predicted by the model also have masses at this scale, but the mass of the third is closer to 103 GeV, the experimental lower limit for an additional charged lepton [22]. An effective theory below the scale of $\tilde{M}_{\nu}$ that excluded those heavy leptons would have an anomaly (see [24, 25]).
Applying arguments from [26], this effective anomaly implies the following upper limit for the mass of those neutrinos:
$\tilde{M}_{\nu}\lesssim\frac{64\pi^{3}M_{Z}}{|3+2-2*2^{3}||\tfrac{1}{2}\sqrt{\tfrac{3}{5}}g_{Y}|^{3}}\simeq 32\textrm{ TeV}.$ (4.1)
When $\tilde{M}_{\nu}$ obeys the above relation, the $Z$ boson mass is able to resolve the obstruction to renormalizability of the effective theory that is generated by anomalous diagrams involving the $Y$ boson.
This upper limit depends on how many of the charged Omega leptons have masses similar to $\tilde{M}_{\nu}$. If all three of them do, then the limit is lower ($\sim$19 TeV). If none of them do (they all have masses nearer 103 GeV), then the limit is much higher since the effective anomaly is not as severe ($\sim$120 TeV). If only one has mass similar to $\tilde{M}_{\nu}$, then the upper limit is $\sim$ 90 TeV.
After the $Z$ and $W$ bosons acquire masses at the electroweak scale, the remaining massless bosons are gluons, photons, and the $Z^{\prime}$ boson. None of these couple to the heavy neutrinos, so they no longer generate an effective anomaly below the electroweak scale. Also, the right-handed Omega leptons ($\Omega_{R}$) and the Hermitian conjugates of the left-handed Omega leptons ($\bar{\Omega}_{L}$) have equal and opposite charges for the photon and $Z^{\prime}$, so the Omega leptons do not generate an effective anomaly below the electroweak scale. But another effect causes an effective anomaly for the $Z^{\prime}$ boson.
The different right-handed components of the top quark ($t_{R}$ and $\bar{t}_{L}$) have opposite electric charges, but they do not have opposite $Z^{\prime}$ charges. Therefore, an effective theory without the top quark has an anomaly for the $Z^{\prime}$ boson. Since the $Z^{\prime}$ boson does not acquire mass at the electroweak scale, it is proposed that this effective anomaly drives $g_{Z^{\prime}}$ to a very small value.
Said another way, since the underlying theory with the top quark is anomaly free, the $Z^{\prime}$-induced anomaly of an effective theory without it must manifest itself by causing the $Z^{\prime}$ coupling to run much more quickly than usual to very small values as the scale under consideration is lowered. In other words, the effective anomaly provides the remainder of the qualitative justification for the very small values of $g_{Z^{\prime}}$ and $\sin\phi_{Z}$ of this model.
At a scale $\mu$ below the unification scale but above $\tilde{\nu}_{M}$, the running coupling constants for the SU(N) groups obey the following equation:
@@
where $n_{f}$ and $n_{s}$, $n_{fA}$ and $n_{sA}$ are the numbers of fundamental fermion and scalar N-tuplets, and the numbers of fermion and scalar adjoint representations, respectively. In all cases, the fermions are 2-component Weyl fermions and the scalars are complex.
In this model, $n_{fA}=0$ and $n_{sA}=1$ for each SU(N) group. The particle content to use between $\Lambda_{U}$ and $\tilde{M}_{\nu}$ is the following:
@@
    Singlet scalars: 6 (2x flavors 1 & 2, 2x adjoint),
where in the notation of this paper, a “W” fermion is part of an isodoublet that interacts with the $W$ boson.
@@
With the above particle content, the beta factors for running of couplings between $\Lambda_{U}$ and $\tilde{M}_{\nu}$ are:
@@
$b_{2}$ $=\left(-\frac{11}{3}2+\frac{1}{3}18+\frac{1}{6}6+\frac{1}{3}2\right)=\frac{1}{3}.$
Between the $\tilde{M}_{\nu}$ scale and the electroweak scale, $b_{3}$ remains the same, but $b_{2}$ is reduced due to decoupling of the three heavy neutrinos, two of the three heavy charged leptons (in the default model) and 1 scalar. For the SU(2) coupling, the 18 fermion doublets are reduced to 15 and the 6 scalar doublets are reduced to 5. So between $\Lambda_{U}$ and $\tilde{M}_{\nu}$, $b_{2}=-5/6$.
The unification scale can be found by starting with the measured values of the SU(2) and SU(3) coupling constants at the scale of $M_{Z}$, using $b_{3}=-\frac{13}{3}$ and $b_{2}=-5/6$ to run the couplings up to the scale of $\tilde{M}_{\nu}$, and then using $b_{3}=-\frac{13}{3}$ and $b_{2}=1/3$ to run them up further until they have the same value. Assuming that $\tilde{M}_{\nu}\simeq$32 TeV (the top of the allowed range from eq (4.1)), then the unification scale is:
$\Lambda_{U}\simeq 6\times 10^{16}\textrm{ GeV}.$ (4.5)
@@
In this model, the beta factor for the weak hypercharge coupling $g_{Y}$ between $\Lambda_{U}$ and $\tilde{M}_{\nu}$ is:

### v9

@@ 3.3 Quark masses and mixing
Most of the off-diagonal FCNC elements of these matrices are small, but a few are more significant. The only Z mixing matrix elements with magnitudes greater than 0.003 are $V_{Zs\bar{d}}^{A},V_{Zf\bar{b}}^{L},V_{Zf\bar{s}}^{L},V_{Zb\bar{d}}^{L}$, and $V_{Zt\bar{c}}^{R}$. More detailed work would be needed to verify that the $Z$ boson FCNC mixing of this model is fully consistent with experimental data.
3.4 Neutrinos
In addition to three left-handed neutrinos $\nu_{W}^{{\dagger}(F)}$, this model also has three right-handed heavy neutrinos $\tilde{\nu}_{W}^{(F)}$. From eq (A.13) of the Appendix, quantum effects generate a mass matrix of the form
$\tfrac{1}{3}M_{\nu}^{(2)}\left(\begin{array}[]{ccc}{1}&{1}&{1}\\ {1}&{1}&{1}\\ {1}&{1}&{1}\end{array}\right)\,$, (3.26)
where $M_{\nu}^{(2)}$ is a very large mass, and both the rows and columns are $\tilde{\nu}_{W}^{(1)},\tilde{\nu}_{W}^{(2)},\tilde{\nu}_{W}^{(3)}$. A flavor basis rotation can diagonalize this matrix to
$\left(\begin{array}[]{ccc}{0}&{0}&{0}\\ {0}&{M_{\nu}^{(2)}}&{0}\\ {0}&{0}&{0}\end{array}\right)\,.$ (3.27)
As described in eq (4.5), the upper limit for this mass is $\sim$5200 TeV.
A flavor rotation that accomplishes this diagonalization is given by:
$U_{\nu}=\left(\begin{array}[]{ccc}{\tfrac{2}{\sqrt{6}}}&{\tfrac{1}{\sqrt{3}}}&{0}\\ {\tfrac{-1}{\sqrt{6}}}&{\tfrac{1}{\sqrt{3}}}&{-\tfrac{1}{\sqrt{2}}}\\ {\tfrac{-1}{\sqrt{6}}}&{\tfrac{1}{\sqrt{3}}}&{\tfrac{1}{\sqrt{2}}}\end{array}\right)\left(\begin{array}[]{ccc}{\cos\theta_{\nu}}&{0}&{\sin\theta_{\nu}}\\ {0}&{1}&{0}\\ {-\sin\theta_{\nu}}&{0}&{\cos\theta_{\nu}}\end{array}\right)$ (3.28)
for any value of $\theta_{\nu}$.
At a lower scale in an effective theory that excludes the heaviest neutrino, the same quantum mechanism generates the following new mass term in the rotated basis:
$\left(\begin{array}[]{ccc}{\tfrac{1}{2}M_{\nu}^{(1)}}&{0}&{\tfrac{1}{2}M_{\nu}^{(1)}}\\ {0}&{M_{\nu}^{(2)}}&{0}\\ {\tfrac{1}{2}M_{\nu}^{(1)}}&{0}&{\tfrac{1}{2}M_{\nu}^{(1)}}\end{array}\right)\,.$ (3.29)
Diagonalization of this can be accomplished via a modified choice of $\theta_{\nu}$. At an even lower scale, the quantum mechanism generates the third heavy neutrino mass $M_{\nu}^{(3)}$. The diagonalized heavy neutrino mass matrix is then:
$\left(\begin{array}[]{ccc}{M_{\nu}^{(1)}}&{0}&{0}\\ {0}&{M_{\nu}^{(2)}}&{0}\\ {0}&{0}&{M_{\nu}^{(3)}}\end{array}\right)\,$, (3.30)
where $M_{\nu}^{(2)}>M_{\nu}^{(1)}>M_{\nu}^{(3)}$.
Since $\theta_{\nu}$ can be anything, it is valid to make the choice $\sin\theta_{\nu}\simeq 1/\sqrt{48}$ for the full rotation. In that case
$U_{\nu}\simeq\left(\begin{array}[]{ccc}{0.81}&{0.58}&{0.12}\\ {-0.30}&{0.58}&{-0.76}\\ {-0.51}&{0.58}&{0.64}\end{array}\right)\,.$ (3.31)
[…]
$U_{\rm PMNS}=\left(\begin{array}[]{ccc}{0.82}&{0.55}&{0.15}\\ {-0.27}&{0.61}&{-0.75}\\ {-0.50}&{0.58}&{0.65}\end{array}\right)\,.$ (3.32)
The similarity of $U_{\nu}$ to $U_{\rm PMNS}$ is discussed below.
From eq (A.14) in the Appendix, quantum effects generate additional Dirac-mass terms that connect right- and left-handed neutrinos as follows:
$\left(\begin{array}[]{c}{\tilde{\nu}_{W}^{(1)}}\\ {\tilde{\nu}_{W}^{(2)}}\\ {\tilde{\nu}_{W}^{(3)}}\end{array}\right)^{T}\tfrac{1}{9}m_{\nu}^{\prime}\left(\begin{array}[]{ccc}{-5}&{4}&{4}\\ {4}&{-5}&{4}\\ {4}&{4}&{-5}\end{array}\right)\left(\begin{array}[]{c}{\nu_{W}^{(1)}}\\ {\nu_{W}^{(2)}}\\ {\nu_{W}^{(3)}}\end{array}\right)\,.$ (3.33)
To be consistent with the heavy neutrino basis of eq (3.30), the $\tilde{\nu}_{W}^{(F)}$ neutrinos of eq (3.33) must be rotated to that basis as follows:
$\tilde{\nu}_{W}^{(F)}=\tilde{\nu}_{W}^{\prime(F^{\prime})}(U_{\nu}^{\dagger})_{F^{\prime}F}\,.$ (3.34)
If the left-handed neutrinos experience the same rotation,
$\nu_{W}^{(F)}=(U_{\nu})_{FF^{\prime}}\nu_{W}^{\prime(F^{\prime})}\,$, (3.35)
then the Dirac-mass matrix is diagonalized to
$\left(\begin{array}[]{c}{\tilde{\nu}_{W}^{\prime(1)}}\\ {\tilde{\nu}_{W}^{\prime(2)}}\\ {\tilde{\nu}_{W}^{\prime(3)}}\end{array}\right)^{T}m_{\nu}^{\prime}\left(\begin{array}[]{ccc}{-1}&{0}&{0}\\ {0}&{\tfrac{1}{3}}&{0}\\ {0}&{0}&{-1}\end{array}\right)\left(\begin{array}[]{c}{\nu_{W}^{\prime(1)}}\\ {\nu_{W}^{\prime(2)}}\\ {\nu_{W}^{\prime(3)}}\end{array}\right)\,.$ (3.36)
In this diagonal basis, each flavor of neutrino forms its own seesaw to generate a small left-handed neutrino mass. The magnitudes of these masses are
$m_{\nu}^{(1)}=\frac{(m_{\nu}^{\prime})^{2}}{M_{\nu}^{(1)}}\hskip 14.22636ptm_{\nu}^{(2)}=\frac{(\tfrac{1}{3}m_{\nu}^{\prime})^{2}}{M_{\nu}^{(2)}}\hskip 14.22636ptm_{\nu}^{(3)}=\frac{(m_{\nu}^{\prime})^{2}}{M_{\nu}^{(3)}}\,.$ (3.37)
To reproduce observations, the quantum-generated masses $m_{\nu}^{\prime}$ and $M_{\nu}^{(F)}$ would have to have values that produce the correct differences between masses squared of the above left-handed neutrinos [22]. As one example, in a variant of normal neutrino mass ordering where $(m_{\nu}^{(1)})^{2}-(m_{\nu}^{(2)})^{2}=7.42\times 10^{-5}{\rm eV}^{2}$, the correct mass differences are generated by $m_{\nu}^{\prime}\simeq$ 1.0 MeV, $M_{\nu}^{(2)}\simeq$ 1800 TeV, $M_{\nu}^{(1)}\simeq$ 120 TeV, and $M_{\nu}^{(3)}\simeq$ 20 TeV. For another example, the correct differences are also generated by $m_{\nu}^{\prime}\simeq$ 0.1 MeV, $M_{\nu}^{(2)}\simeq$ 100 TeV, $M_{\nu}^{(1)}\simeq$ 1.2 TeV, and $M_{\nu}^{(3)}\simeq$ 200 GeV.
From eq (3.35), $U_{\nu}$ rotates from the left-handed neutrino mass basis to the flavor basis of this paper. In the next section, it will be argued that the flavor basis of this paper is similar to the charged lepton mass basis. $U_{\rm PMNS}$ is the experimentally measured rotation from the left-handed neutrino mass basis to the charged lepton mass basis. Since $U_{\nu}$ in this model is similar to $U_{\rm PMNS}$, it is suggested that the mechanisms presented here are a good first approximation for the origin of the observed PMNS rotation of the neutrino flavor basis.
3.5 Charged Leptons
@@
From eq (2.18), the last three diagonal components of the photon field when acting on leptons are $e\left(0,1,1\right)$. As a result, if the left- and right-chiral components of the three flavors of normal light charged leptons ($l$) are mapped to $l_{L}^{(F)}=e_{W}^{{\dagger}(F)}$ and $l_{R}^{(F)}=\tilde{e}^{(F)}$, then those leptons have the correct electric charge of 1-1 as well as the correct isodoublet and isosinglet designations. These light leptons get their masses from seesaw mechanisms together with this model’s additional heavy charged leptons discussed below.
This model has three flavors of heavy charged leptons that are called “Omega” leptons. The three flavors of heavy and light leptons ($\Omega$ and $l$) are identified as follows:
$\begin{array}[]{ccc}{\nu_{W}^{(F)}=\nu_{l}^{{\dagger}(F)}}&{e_{W}^{(F)}=l_{L}^{{\dagger}(F)}}&{e^{(F)}=\Omega_{L}^{{\dagger}(F)}}\\ {\tilde{\nu}_{W}^{(F)}=\nu_{\Omega}^{(F)}}&{\tilde{e}_{W}^{(F)}=\Omega_{R}^{(F)}}&{\tilde{e}^{(F)}=l_{R}^{(F)},}\end{array}$ (3.38)
@@
From this identification, it can be seen that the negatively charged light leptons have the correct behavior that only their left-handed components interact with the $W$ boson, forming isodoublets with their neutrinos. On the other hand, the opposite is true for negatively charged Omega leptons; their $W$ boson interactions are right-handed. The light leptons not only have the same interactions with the photon and $W$ boson as in the Standard Model, they also have the same interactions with the $Z$ boson, as can be seen from eq (2.20).
This model’s charged leptons have a $6\times 6$ mass matrix of the form:
$\left(\begin{array}[]{c}{\Omega_{R}^{(F)}}\\ {l_{R}^{(F)}}\end{array}\right)^{T}\left(\begin{array}[]{cc}{\bar{\varphi}_{2}\tilde{\Gamma}_{1FF^{\prime}}}&{\tilde{m}_{WFF^{\prime}}^{\prime}}\\ {\tilde{m}_{SFF^{\prime}}^{\prime}}&{0}\end{array}\right)\left(\begin{array}[]{c}{\Omega_{L}^{{\dagger}(F^{\prime})}}\\ {l_{L}^{{\dagger}(F^{\prime})}}\end{array}\right)\,.$ (3.39)
The couplings and masses in the above matrix could have both tree-level and quantum-generated contributions. As shown in eq (A.15) of the appendix, the quantum-generated contribution to $\tilde{m}_{WFF^{\prime}}^{\prime}$ is diagonal in flavor (only nonzero for $F^{\prime}=F$). To a first approximation, the whole matrix will be assumed to be diagonal in flavor, becoming three separate seesaw matrices for the three flavors of charged leptons.
In this approximation, the masses of the three flavors of heavy and light leptons are given by:
$m_{\Omega}^{(F)}\simeq\bar{\varphi}_{2}\tilde{\Gamma}_{1FF}\hskip 14.22636ptm_{l}^{(F)}\simeq\frac{\tilde{m}_{WFF}^{\prime}\tilde{m}_{SFF}^{\prime}}{\bar{\varphi}_{2}\tilde{\Gamma}_{1FF}}\,.$ (3.40)
As discussed in the anomaly part of section 4, the masses of the Omega leptons should be in similar ranges to those of their partner right-handed neutrinos. If they were equal to the second neutrino mass example above, the Omega lepton masses could be $m_{\Omega}^{(1)}\simeq$ 100 TeV, $m_{\Omega}^{(2)}\simeq$ 1.2 TeV, and $m_{\Omega}^{(3)}\simeq$ 200 GeV. One requirement is that the Omega lepton masses must be larger than the lower limit of 103 GeV established in [23].
Two of the lepton gauge eigenstates from eq (3.39) can be expressed in terms of lepton mass eigenstates as follows:
${\Omega_{R}^{(F)}}$ $\simeq\Omega_{mR}^{(F)}+(\tilde{m}_{SFF}^{\prime}/\bar{\varphi}_{2}\tilde{\Gamma}_{1FF})l_{mR}^{(F)}$ (3.41)
${\Omega_{L}^{{\dagger}(F)}}$ $\simeq\Omega_{mL}^{{\dagger}(F)}-(\tilde{m}_{WFF}^{\prime}/\bar{\varphi}_{2}\tilde{\Gamma}_{1FF})l_{mL}^{{\dagger}(F)}\,$,
where a subscript “mm” denotes a mass eigenstate.
By defining the propagating part $h_{b}$ of the scalar $(\varphi_{2})_{32}$ as follows
${\rm Im}((\varphi_{2})_{32})=\bar{\varphi}_{2}/\sqrt{2}+\tfrac{1}{2}h_{b}\,$, (3.42)
it can be seen that eq (3.39) also implies an interaction between the scalar $h_{b}$ and the lepton states $\Omega_{R}^{(F)}$ and $\Omega_{L}^{{\dagger}(F)}$. Then using eq (3.41), it can be seen that this model predicts $h_{b}$ decay to the standard light lepton mass eigenstates:
$h_{b}\to l_{mL}^{{\dagger}(F)}l_{mR}^{(F)}\propto\frac{\tilde{m}_{WFF}^{\prime}\tilde{m}_{SFF}^{\prime}}{\bar{\varphi}_{2}^{2}\tilde{\Gamma}_{1FF}}\simeq\frac{m_{l}^{(F)}}{\bar{\varphi}_{2}}\simeq\frac{m_{l}^{(F)}}{m_{b}}\,$, (3.43)
where eq (3.40) was used.
@@ 4 Unification and Anomalies
As mentioned in section 1, at a scale above any symmetry breaking, the Twisted Superfields theory is free of anomalies. The reason is simple: for every fermion in every representation, there is another fermion in a conjugate representation of each of the SU(3) groups that also has opposite charges for the Abelian fields $A_{+\mu}^{0}$ and $A_{-\mu}^{0}$. Consequently, all of the chiral gauge anomaly triangle diagrams cancel – just as they do for SQCD.
Another way to describe anomaly cancellation is the following: Because of the theory’s fundamental-conjugate structure, it is possible to pair up 2-component fermions into Dirac fermions such that all fermion gauge interactions are vector currents (no axial vector currents). For example, vector-current $W$ boson interactions connect ($c_{L}$, $c_{R}$) and ($s_{L}$, $f_{R}$) as well as ($t_{L}$, $u_{R}$) and ($b_{L}$, $d_{R}$). In a model with only vector currents, there are no chiral gauge anomalies.
Below certain energies, mass eigenstates of fermions become better descriptions than the above-described Dirac pairings. When fermions are put into mass eigenstates, the theory has axial-vector currents connecting them. In other words, axial-vector currents (and gauge anomalies) are not a structural part of the theory; they are just an effect of mass at low energies.
[…]
Below the unification scale $\Lambda_{U}$, the next lower scale is defined by the masses $M_{\nu}^{(2)}$ and $m_{\Omega}^{(1)}$ of the heaviest right-handed neutrino and charged Omega lepton that were discussed in the last section. At a scale $\mu$ between $\Lambda_{U}$ and $M_{\nu}^{(2)}$, the running coupling constants for the SU(N) groups obey the following equation:
@@
where $n_{f}$ and $n_{s}$, $n_{fA}$ and $n_{sA}$ are the numbers of fundamental fermion and scalar N-tuplets, and the numbers of fermion and scalar adjoint representations, respectively. In all cases, the fermions are 2-component fermions and the scalars are complex.
In this model, $n_{fA}=0$ and $n_{sA}=1$ for each SU(N) group. The particle content to use between $\Lambda_{U}$ and $M_{\nu}^{(2)}$ is the following:
@@
    Singlet scalars: 6 (2x flavors 1 & 2, 2x adjoint) .
In this notation, triplet means an SU(3) color triplet, and a “W” fermion is part of an isodoublet that interacts with the $W$ boson.
@@
With the above particle content, the beta factors for running of couplings between $\Lambda_{U}$ and $M_{\nu}^{(2)}$ are:
@@
$b_{2}$ $=\left(-\frac{11}{3}2+\frac{1}{3}12+\frac{1}{6}6+\frac{1}{3}2\right)=\frac{1}{3}.$
The last section presented the assumption that each of the heavy neutrino and Omega lepton masses fall into similar ranges. Due to this assumption, in an effective theory that excludes the heaviest of each, all anomalous triangle diagrams involving only SU(2) fields cancel. However, triangle diagrams involving the Abelian $Y$ and/or $Y^{\prime}$ fields do not cancel; they generate an anomaly in the effective theory [24, 25].
This situation is the one that was studied in [26]: An anomaly-free theory exists, but below a certain scale, an effective theory that excludes some physics has an anomaly. This anomaly gets resolved by gauge bosons acquiring masses. Turning it around, given an anomalous effective theory with massive gauge bosons, the paper calculated the upper limit for the scale at which new physics must appear. Those arguments can be used in this theory to calculate an upper limit for the heaviest lepton masses.
In an effective theory without the heaviest neutrino and Omega lepton, two “mixed” anomalies are generated. These come from diagrams involving two SU(2) fields and either a $Y$ or a $Y^{\prime}$. From eqs (2.14) and (2.15), the mixed anomaly traces for these leptons are:
${\cal A}_{Y}$ $={\rm Tr}\left(T_{2}^{3}T_{2}^{3}T_{21}^{Y}\right)=\tfrac{1}{4}\sqrt{\tfrac{3}{5}}$ (4.4)
${\cal A}_{Y^{\prime}}$ $={\rm Tr}\left(T_{2}^{3}T_{2}^{3}T_{21}^{Y^{\prime}}\right)=\tfrac{1}{2\sqrt{10}}\,.$
From eq (4.13) of [26], the upper limit on the mass of the heaviest leptons is:
[…]
$M_{\nu}^{(2)}\sim m_{\Omega}^{(1)}$ $\lesssim\frac{64\pi^{3}M_{Z^{\prime}}}{|{\cal A}_{Y^{\prime}}g_{Y^{\prime}}|g_{2}^{2}}\,$,
where coupling constants are evaluated at the $Z$ and $Z^{\prime}$ boson mass scales, respectively. Anomalies are also generated by “non-mixed” diagrams having only $Y$ and/or $Y^{\prime}$ gauge fields. These anomalies lead to larger upper limits, so will be ignored.
If $M_{\nu}^{(2)}$ is equal to the same percentage of both of the mixed upper limits, the following relation holds:
$g_{Y^{\prime}}=\sqrt{\tfrac{3}{2}}\frac{M_{Z^{\prime}}}{M_{Z}}g_{Y}\,$, (4.6)
where $M_{Z^{\prime}}$ is the mass of the model’s $Z^{\prime}$ boson. In section 5, it is suggested that a good candidate for this model’s $Z^{\prime}$ boson is the X17 boson with a mass of 17 MeV. If that were the case, then it would imply that $g_{Y^{\prime}}\simeq 3\times 10^{-4}g_{Y}$, the correct scale to reproduce X17 data. In this model, it is assumed that in order to satisfy eq (4.5), the effective anomaly drives the $Y^{\prime}$ coupling down to a much smaller value than would be expected from running of the coupling constant in an anomaly-free effective theory.
In order to estimate the unification scale, the simplifying assumption is made that the model’s leptoquarks have masses similar to those of the heaviest leptons. As a result, between that heavy mass scale ($M_{\nu}^{(2)}$) and the electroweak scale, $b_{3}$ is reduced to 16/3-16/3 due to decoupling of the leptoquarks. In that same range, $b_{2}$ is also reduced due to decoupling of the heaviest leptons and the one very heavy scalar mentioned in eq (A.13) of the appendix. In other words, for the SU(2) coupling, the 18 fermion doublets are reduced to 17 and the 6 scalar doublets are reduced to 5, resulting in $b_{2}=-1/6$.
The unification scale can be found by starting with the measured values of the SU(2) and SU(3) coupling constants at the scale of $M_{Z}$, using $b_{3}=-16/3$ and $b_{2}=-1/6$ to run the couplings up to the scale of $M_{\nu}^{(2)}$, and then using $b_{3}=-13/3$ and $b_{2}=1/3$ to run them up further until they have the same value. Using $M_{\nu}^{(2)}\simeq$100 TeV, the unification scale is:
$\Lambda_{U}\simeq 10^{14}\textrm{ GeV}.$ (4.7)
@@
In this model, the beta factor for the weak hypercharge coupling $g_{Y}$ between $\Lambda_{U}$ and $M_{\nu}^{(2)}$ is: